### v1

# Partial comonotonicity and distortion riskmetrics

Abstract
We establish a connection between subclasses of distortion riskmetrics and dependence structures, ensuring their additivity. A new notion of positive dependence, called partial comonotonicity, is developed, which nests the existing concepts of comonotonicity and single-point concentration. For two random variables, being comonotonic with a third one does not imply that they are comonotonic; instead, this defines an instance of partial comonotonicity. Any specific instance of partial comonotonicity uniquely characterizes a class of distortion riskmetrics through additivity under this dependence structure. An implication of this result is the characterization of the Expected Shortfall using single-point concentration.
Keywords: Partial comonotonicity, distortion risk measures, positive dependence, additivity
1 Introduction
Positive dependence among risks is an important consideration in risk management (McNeil et al. (2015)). As the extreme form of positive dependence, comonotonicity models perfect comovement between random losses from a financial portfolio. Additivity for comonotonic losses is a key property in the characterization of many risk measures and functionals, widely used in economics, finance and insurance (Schmeidler (1989); Wang et al. (1997); Kusuoka (2001)). A weaker notion of positive dependence, known as $p$-concentration, models the following scenario: there is some catastrophic event with probability $(1-p)$, under which all assets in a portfolio perform worse than they would have if such an event had not occurred. In a characterization of the Expected Shortfall at level $p$ (Wang and Zitikis (2021)), $p$-concentration played a central role. The current paper aims to characterize subclasses of distortion risk measures, and more generally, distortion riskmetrics (not requiring the distortion function to be increasing, defined in Wang et al. (2020a)), via their additivity for specific dependence structures. We achieve the characterization by using a new collection of positive dependence structures called partial comonotonicity.
Notably, $p$-concentration, upper comonotonicity (Cheung (2009)) and comonotonicity are particular cases of partial comonotonicity. A broader generalization called weak comonotonicity was introduced in (Wang and Zitikis (2020)). There, random variables are required to move in the same direction on average in a weighted sense. By choosing a large number of simple weights, partial comonotonicity can be recovered.
A dependence structure $D$ in this paper may refer to a copula or a particular instance of partial comonotonicity, i.e., partial comonotonicity indexed or parametrized by some mathematical object. We will see that a particular instance $D$ of partial comonotonicity corresponds to a collection $\mathcal{C}_{D}$ of copulas. That is, a random vector has the dependence structure $D$ (from the class of partial comonotonicity) if and only if its copula is in $\mathcal{C}_{D}$. For a dependence structure $D$, a riskmetric $f$ is called $D$-additive if for any random variables $(X,Y)$ satisfying $D$, $f$ is additive for $(X,Y)$. When $D$ is chosen as comonotonicity, we arrive at the usual comonotonic additivity, the defining property of Choquet integrals (Schmeidler (1986)). The law-invariant case was characterized by Yaari (1989). In this case, the Choquet integrals are also known as distortion risk measures, studied by Dhaene et al. (2002, 2006, 2012), and their non-monotone version is studied by Wang et al. (2020b). On the other hand, $p$-concentration-additivity ($p$-additivity for short) is a characterizing property of the Expected Shortfall at level $p$, as mentioned above. Using the desired additivity properties to narrow down the potential choices of risk measures is a recurring theme in the study of risk measures; see Mu et al. (2024); Bellini et al. (2021); Principi et al. (2024) for some recent developments. We follow this classic approach in this paper. Since any instance of partial comonotonicity is weaker than comonotonicity and the corresponding additivity properties are reversed in strength, any functional satisfying any instance of partial comonotonic additivity must be comonotonic additive. When studying the aggregation of risk, we focus on the dependence structure of the random losses, and hence, law invariance is also a standard assumption. Therefore, we will be working with distortion risk measures and their non-monotone generalization, distortion riskmetrics.
In Section 3, we formulate partial comonotonicity by using subsets $K$ of $[0,1]$. From this formulation, partial comonotonicity can be seen as a natural generalization of (single-point) $p$-concentration; thus, we call it $K$-concentration. Section 4 develops an equivalent formulation of partial comonotonicity through a parametrization by left-continuous increasing functions $g:[0,1]\mapsto\mathbb{R}$. This formulation enables us to obtain an elegant result in Section 5: for any random vector, a spectral risk measure with risk spectrum $g$ is additive precisely when partial comonotonicity, parametrized by $g$, holds.
2 Notations, conventions, and some basic facts
Throughout, increasing functions are in the weak sense, and inequalities among vectors are componentwise. We typically use capital letters, such as $X$, to denote random variables and bold capital letters, such as $\mathbf{X}$, to denote random vectors.
Let $\mathcal{X}=L^{\infty}$ be the collection of essentially bounded random variables on an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, with almost surely equal objects treated as identical. We refer to the collection of functionals $\rho:\mathcal{X}\mapsto(-\infty,\infty]$ as riskmetrics. Commonly used riskmetrics include the Value-at-Risk (VaR) at level $p\in(0,1)$:
$\mathrm{VaR}_{p}(X)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leq x)\geq p\},$
and the Expected Shortfall (ES) at level $p\in(0,1)$:
$\mathrm{ES}_{p}(X)=\frac{1}{1-p}\int_{p}^{1}\mathrm{VaR}_{q}(X)\mathrm{~{}d}q.$
We also define $\mathrm{VaR}_{1}(X)$ and $\mathrm{ES}_{1}(X)$ to be $\mathrm{ess\text{-}sup}(X)$, the essential supreright-continuousa fixed random variable $X$, the left quantile function $Q^{-}_{X}:(0,1]\mapsto\mathbb{R}$, often written as $Q_{X}$, is defined as $Q_{X}(p)=\mathrm{VaR}_{p}(X)$ and the right quantile function of $X$, $Q_{X}^{+}:[0,1)\mapsto\mathbb{R}$ is defined as
$Q^{+}_{X}(p)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leq x)>p\}.$
We write the corresponding right-continuous Value-at-Risk as $\mathrm{VaR}^{+}_{p}(X)$. Note that at $p=0$, $Q_{X}^{+}(0)$ is the essential infimum. A coherent risk measure (Artzner et al. (1999)) is a riskmetric $\rho$ satisfying the following properties:
  • (M) Monotonicity: $\rho(X)\leq\rho(Y)$ for all $X,Y\in\mathcal{X}$ with $X\leq Y$,
  • (TI) Translation invariance: $\rho(X+c)=\rho(X)+c$ for all $X\in\mathcal{X}$ and $c\in\mathbb{R}$,
  • (SA) Subadditivity: $\rho(X+Y)\leq\rho(X)+\rho(Y)$ for all $X,Y\in\mathcal{X}$,
  • (PH) Positive homogeneity: $\rho(\lambda X)=\lambda\rho(X)$ for all $X,Y\in\mathcal{X}$ and $\lambda\geq 0$,
and condition (SA) may be substituted with the following:
  • (CX) Convexity: $\rho(\alpha X+(1-\alpha)Y)\leq\alpha\rho(X)+(1-\alpha)\rho(Y)$ for all $X,Y\in\mathcal{X}$ and $\alpha\in(0,1)$.
A riskmetric satisfying (M) is referred to as a risk measure. If, in addition, (TI) is satisfied, the risk measure is monetary. These two properties allow us to interpret a monetary risk measure as a capital requirement. A random vector $(X_{1},\dots,X_{d})\in\mathcal{X}^{d}$ is called comonotonic if there exists a random variable $Z$ and increasing functions $f_{1},\dots,f_{d}:\mathbb{R}\mapsto\mathbb{R}$ such that
$(X_{1},\dots,X_{d})=\left(f_{1}(Z),\dots,f_{d}(Z)\right).$
Two more commonly seen properties of riskmetrics are
  • (LI) Law invariance: $\rho(X)=\rho(Y)$ whenever $X\buildrel\mathrm{d}\over{=}Y$.
  • (CA) Comonotonic additivity: $\rho(X_{1})+\dots+\rho(X_{d})=\rho(X_{1}+\dots+X_{d})$ whenever $(X_{1},\dots,X_{d})$ is comonotonic.
A spectral risk measure $\rho$ as introduced in Acerbi (2002) can be represented in the form
$\rho(X)=\rho_{g}(X)=\int_{0}^{1}g(t)Q_{X}(t)\mathrm{~{}d}t$
where $g:[0,1]\mapsto\mathbb{R}$ is non-negative, increasing and integrates to $1$. Such a function $g$ is referred to as a risk spectrum. Spectral risk measures are coherent and satisfy (LI) and (CA). Kusuoka (2001) further showed that all coherent risk measures that satisfy (LI) and (CA) are of the form
$\alpha\rho_{g}+(1-\alpha)\mathrm{ess\text{-}sup}$
for some spectral risk measure $\rho_{g}$ and $\alpha\in[0,1]$. A distortion riskmetric is a riskmetric of the form
$I_{h}(X)=\int_{-\infty}^{0}(h(\mathbb{P}(X>x))-h(1)\mathrm{~{}d}x+\int_{0}^{\infty}h(\mathbb{P}(X>x))\mathrm{~{}d}x$
where $h:[0,1]\mapsto\mathbb{R}$, called the distortion function, is of bounded variation with $h(0)=0$. We denote by $\mathcal{H}^{\rm bv}$ the set of all distortion functions. The functional $I_{h}$ is also known as a signed Choquet integral. Such functionals are characterized in Wang et al. (2020b) as riskmetrics satisfying (LI) and (CA) together with the following continuity condition:
  • (UNC) Uniform norm-continuity: $\rho$ is uniformly continuous with respect to $L^{\infty}$-norm.
The conjugate function of a distortion function $h$ is defined as $\hat{h}(t)=h(1)-h(1-t)$. If $h$ is increasing with $h(1)=1$ and $\hat{h}$ is differentiable with an increasing derivative $g$, we have $I_{h}=\rho_{g}$ so that spectral risk measures are a special case of distortion riskmetrics. The following facts about distortion riskmetrics will be useful.
Fact 1.
Let $h\in\mathcal{H}^{\rm bv}$ and $X\in L^{\infty}$.
  1. (a)
    if $h$ is right-continuous, then $I_{h}(X)=\int_{0}^{1}Q_{X}^{+}(1-p)\mathrm{~{}d}h(p)$;
  2. (b)
    if $h$ is left-continuous, then $I_{h}(X)=\int_{0}^{1}Q_{X}(1-p)\mathrm{~{}d}h(p)$;
  3. (c)
    $I_{h}$ satisfies (PH);
  4. (d)
    $I_{h}$ satisfies (CX) if and only if $h$ is concave.
An event $A\in\mathcal{F}$ is called a tail event for a random variable $X$ if
$X(\omega)\geq X(\omega^{\prime})\mbox{~{}holds a.s. for all }\omega\in A,~{}\omega^{\prime}\in A^{c},$
where $A^{c}$ stands for the complement of $A$. If a tail event $A$ has $\mathbb{P}(A)=1-p$ for $p\in[0,1]$, then $A$ is called a $p$-tail event. A random vector is said to be $p$-concentrated if all components share a common $p$-tail event. We collect the following useful facts for $p$-concentration from Wang and Zitikis (2021).
Fact 2.
The following properties of $p$-concentration hold.
  1. (a)
    A random vector is $q$-concentrated for all $q\in(0,1)$ if and only if it is comonotonic.
  2. (b)
    Suppose that $(X_{1},\dots,X_{d})$ is $p$-concentrated for some $p\in(0,1)$. A set $A$ is a $p$-tail event of $X_{1}+\dots+X_{d}$ if and only if $A$ is a $p$-tail event for each $X_{i}$ for $i=1,\dots,d$;
  3. (c)
    If $(X_{1},\dots,X_{d})$ is $p$-concentrated, then we have the inequalities
    $\mathrm{VaR}_{p}\left(\sum_{i=1}^{d}X_{i}\right)\leq\sum_{i=1}^{d}\mathrm{VaR}_{p}(X_{i})\leq\sum_{i=1}^{d}\mathrm{VaR}^{+}_{p}(X_{i})\leq\mathrm{VaR}^{+}_{p}\left(\sum_{i=1}^{d}X_{i}\right).$
3 Set-indexed partial comonotonicity
The first of the two formulations of partial comonotonicity comes as a generalization of $p$-concentration.
Definition 1.
Given $K\subseteq[0,1]$, we say a random vector $\mathbf{X}$ is $K$-concentrated when $\mathbf{X}$ is $p$-concentrated for each $p\in K$.
Because any random vector $\mathbf{X}$ is trivially $1$-concentrated and 00-concentrated, we sometimes omit 00 and/or $1$ from the set $K$, which does not change the property. Any $K$-concentration is called a dependence structure, but we will also call a copula (Nelson (2006)) a dependence structure.
Example 1.
By choosing the set $K$, we can recover many well-studied dependence structures.
  1. (a)
    For $p\in(0,1)$, $\{p\}$-concentration is precisely $p$-concentration. This is the most distinguishing property used to characterize $\mathrm{ES}_{p}$ in Wang and Zitikis (2021).
  2. (b)
    With $K_{c}=[0,1]$, $K_{c}$-concentration is equivalent to comonotonicity (Fact 2, item (a)). This justifies the term partial comonotonicity since $K$ is always a subset of $[0,1]$.
  3. (c)
    For $p\in(0,1)$, let $K_{p}=[p,1]$. For $\mathbf{X}$ to be $K$-comonotonic, there exists a common $p$-tail event of $\mathbf{X}$ such that the conditional distribution of $\mathbf{X}$ on this event is comonotonic. A similar concept called upper comonotonicity was introduced Cheung (2009), which also models comonotonicity in some tail events.
The next proposition shows that it suffices to consider concentration indexed by closed subsets of $[0,1]$.
Proposition 1.
For $K\subseteq[0,1]$ with its closure denoted by $\mathsf{cl}(K)$, $K$-concentration is equivalent to $\mathsf{cl}(K)$-concentration.
Proof.
If $p_{n}\downarrow p$ with each $p_{n}\in K$, let $A=\bigcup_{n\in\mathbb{N}}A_{p_{n}}$, where $A_{p_{n}}$ is a common $p_{n}$-tail event of $\mathbf{X}$ for each $n$. Since $\mathbb{P}$ satisfies $\sigma$-additivity, we have $\mathbb{P}(A)=1-p$. If $\omega\in A$, then $\omega\in A_{p_{n}}$ for some $n$. Since $A^{c}\subseteq A_{p_{n}}^{c}$, for any $\omega^{\prime}\in A^{c}$ we have $\mathbf{X}(\omega)\geq\mathbf{X}(\omega^{\prime})$, and hence $A$ is a $p$-tail event for $\mathbf{X}$. Similarly, if $p_{n}\uparrow p$ with each $p_{n}\in K$, by continuity from above, $\mathbf{X}$ is $p$-concentrated. ∎
We denote by $\mathcal{S}$ the set of all closed subsets of $[0,1]$. The following proposition proposes a structured approach to selecting the common $p$-tail events for different levels $p\ inK$.
Proposition 2.
Let $K\in\mathcal{S}$ and suppose a random vector $\mathbf{X}$ is $K$-concentrated. Then there exists a collection sets $\{A_{p}\}_{p\in K}$ such that
  1. (a)
    for each $p\in K$, $A_{p}$ is a common $p$-tail event for $\mathbf{X}$;
  2. (b)
    for $p,q\in K$ with $p>q$, we have $A_{p}\subseteq A_{q}$.
Proof.
Let $S=X_{1}+\dots+X_{n}$. By Lemma A.32 of Föllmer and Schied (2016), there exists a uniform $[0,1]$ random variable $U$ such that $Q_{S}(U)=S$ almost surely. For each $p\in K$, let $A_{p}=\{U>p\}$. Clearly, each $A_{p}$ is a $p$-tail event for $S$ and the collection satisfies condition (b). By item (b) of Fact 2, each $A_{p}$ is also a common $p$-tail event for $\mathbf{X}$. ∎
We call $\{A_{p}\}_{p\in K}$ satisfying (i) and (ii) above a nested collection of common tail events.
We are ready to develop results regarding the additivity of quantile functions. We start by introducing the notion of partial comonotonic additivity.
Definition 2.
Let $K\in\mathcal{S}$. A riskmetric $\rho:\mathcal{X}\mapsto\mathbb{R}$ is $K$-additive if for all $K$-concentrated random vectors $\mathbf{X}=(X_{1},\dots,X_{d})$, we have
$\rho\left(\sum_{i=1}^{d}X_{i}\right)=\sum_{i=1}^{d}\rho(X_{i}).$
Theorem 1.
Let $p\in(0,1]$ and $K$ be a closed subset of $[0,1]$. The following are equivalent:
  1. (a)
    $\mathrm{VaR}_{p}$ is $K$-additive.
  2. (b)
    There exists a strictly increasing sequence $p_{n}\uparrow p$ with each $p_{n}\in K$.
    Similarly, for $p\in[0,1)$, the following are equivalent:
  3. (c)
    $\mathrm{VaR}^{+}_{p}$ is $K$-additive.
  4. (d)
    There exists a strictly decreasing sequence $p_{n}\downarrow p$ with each $p_{n}\in K$.
Proof.
Let $p_{n}\uparrow p$ and $\mathbf{X}=(X_{1},\dots,X_{d})$ be $p_{n}$-concentrated for each $n\in\mathbb{N}$. By (c) from Fact 2, for each $n\in\mathbb{N}$ we have the inequalities
$\sum_{i=1}^{d}\mathrm{VaR}_{p_{n}}(X_{i})\leq\sum_{i=1}^{d}\mathrm{VaR}^{+}_{p_{n}}(X_{i})$ $\leq\mathrm{VaR}^{+}_{p_{n}}\left(\sum_{i=1}^{d}X_{i}\right)$
$\leq\mathrm{VaR}_{p_{n+1}}\left(\sum_{i=1}^{d}X_{i}\right)\leq\sum_{i=1}^{d}\mathrm{VaR}_{p_{n+1}}(X_{i}).$
As $n\to\infty$, by the left-continuity of the left quantile function, the first and last terms converge to $\sum_{i=1}^{d}{\mathrm{VaR}_{p}(X_{i})}$ and the last two terms are equal in the limit.
Similarly for $p_{n}\downarrow p$, we have
$\sum_{i=1}^{d}\mathrm{VaR}^{+}_{p_{n}}(X_{i})\geq\sum_{i=1}^{d}\mathrm{VaR}_{p_{n}}(X_{i})$ $\geq\mathrm{VaR}_{p_{n}}\left(\sum_{i=1}^{d}X_{i}\right)$
$\geq\mathrm{VaR}^{+}_{p_{n+1}}\left(\sum_{i=1}^{d}X_{i}\right)\geq\sum_{i=1}^{d}\mathrm{VaR}^{+}_{p_{n+1}}(X_{i}).$
As $n\to\infty$, by the right-continuity of the right quantile function, the first and last terms converge to $\sum_{i=1}^{d}{\mathrm{VaR}^{+}_{p}(X_{i})}$ and the last two terms are equal in the limit. ∎
The following proposition enables us to focus on the special case $d=2$ for many of the proofs that follow.
Proposition 3.
Fix $d\in\mathbb{N}$, $K\in\mathcal{S}$ and $\mathbf{X}\in\mathcal{X}^{d}$. We have that $\mathbf{X}$ is $K$-concentrated if and only if for any $k\in\mathbb{N}$ and any (componentwise) increasing function $\mathbf{f}=(f_{1},\dots,f_{k}):\mathbb{R}^{d}\mapsto\mathbb{R}^{k}$, $\mathbf{f}(\mathbf{X})$ is $K$-concentrated.
Proof.
Take $p\in K$ and let $A_{p}$ be a common $p$-tail event for $\mathbf{X}$. For a.s. all $\omega\in A_{p}$ and $\omega^{\prime}\in A_{p}^{c}$, we have $\mathbf{X}(\omega)\geq\mathbf{X}(\omega^{\prime})$. Since $\mathbf{f}$ is increasing, $[\mathbf{f}(\mathbf{X})](\omega)\geq[\mathbf{f}(\mathbf{X})](\omega^{\prime})$ so that $A_{p}$ is a common $p$-tail event for $\mathbf{f}(\mathbf{X})$, and hence $\mathbf{f}(\mathbf{x})$ is $p$-concentrated. ∎
The invariance property under increasing transforms in Proposition 3 is also satisfied by the popular dependence concept of positive association, introduced by Esary et al. (1967), who showed that if $\mathbf{X}$ is positively associated, then so is $\mathbf{f}(\mathbf{X})$ for any increasing function $\mathbf{f}$. Proposition 3 generalizes the fact that comonotonicity satisfies this invariance, shown by Lauzier et al. (2023).
Remark 1.
Suppose $\rho$ is additive for $2$-dimensional $K$-concentrated vectors. That is, $\rho(X+Y)=\rho(X)+\rho(Y)$ for any $K$-concentrated random vector $(X,Y)$. Since addition is a pointwise increasing function, for a $K$-concentrated random vector $\mathbf{X}=(X_{1},X_{2},X_{3})$, $(X_{1},X_{2}+X_{3})$ and $(X_{2},X_{3})$ are both $K$-concentrated. So that
$\rho(X_{1}+X_{2}+X_{3})=\rho(X_{1}+(X_{2}+X_{3}))$ $=\rho(X_{1})+\rho(X_{2}+X_{3})$
$=\rho(X_{1})+\rho(X_{2})+\rho(X_{3}).$
This shows that considering the case $d=2$ does not lose generality.
We are ready to develop criteria for $K$-additivity for general distortion riskmetrics.
For any $K\in\mathcal{S}$, we decompose $K^{c}$ as a disjoint union of open intervals $\{B_{i}\}_{i\in\mathbb{N}}$. For a subset $B$ of $[0,1]$, let $1-B=\{1-x:x\in B\}$.
Theorem 2.
For $K\in\mathcal{S}$ and $h\in\mathcal{H}^{\rm bv}$, a distortion riskmetric $I_{h}$ is $K$-additive if and only if $h$ is linear on each $\mathsf{cl}(1-B_{i}$).
Proof.
Given $h$ with bounded variation, decompose $h=h_{l}+h_{r}+h_{c}$ where $\hat{h}_{l}$ (resp. $\hat{h}_{r}$) is constant except for jumps that are left- (resp. right-) continuous, and $h_{c}$ is continuous. Let $\hat{h},\hat{h}_{c},\hat{h}_{l},\hat{h}_{r}$ be corresponding conjugate functions. The functional $I_{h_{l}}$ is a countable weighted sum of left quantiles, and similarly, $I_{h_{r}}$ is a countable weighted sum of right quantiles. Note that $h$ being continuous on $\mathsf{cl}(1-B_{i})$ is equivalent to $\hat{h}$ being continuous on $\mathsf{cl}(B_{i})$.
By Fact 1 (a) and (b), we have
$I_{h}$ $=I_{h_{c}}+I_{h_{l}}+I_{h_{r}}$
$=\int_{0}^{1}Q^{-}_{(~{}\cdot~{})}(p)\mathrm{~{}d}\hat{h}_{c}(p)+\int_{0}^{1}Q^{-}_{(~{}\cdot~{})}(p)\mathrm{~{}d}\hat{h}_{l}(p)+\int_{0}^{1}Q^{+}_{(~{}\cdot~{})}(p)\mathrm{~{}d}\hat{h}_{r}(p).$
Suppose $(X,Y)$ is $K$-concentrated. Let $\{A_{p}\}_{p\in K}$ be a nested family of common tails, we have
$I_{h_{c}}(X+Y)$ $=\int_{0}^{1}Q^{-}_{X+Y}(p)\mathrm{~{}d}\hat{h}_{c}(p)$
$=\sum_{i=1}^{\infty}\int_{K_{i}}Q^{-}_{X+Y}(p)\mathrm{~{}d}\hat{h}_{c}(p)+\sum_{i=1}^{\infty}\int_{B_{i}}Q^{-}_{X+Y}(p)\mathrm{~{}d}\hat{h}_{c}(p).$
Let $m_{i}$ be the slope of $\hat{h}_{c}$ on $B_{i}=(a_{i},b_{i})$, and let $\Omega_{i}=A_{a_{i}}\setminus A_{b_{i}}$, we have:
$\int_{B_{i}}Q^{-}_{X+Y}(p)\mathrm{~{}d}\hat{h}_{c}(p)$ $=m_{i}\mathbb{E}[(X+Y)\mathds{1}_{\Omega_{i}}]$
$=m_{i}\mathbb{E}[X\mathds{1}_{\Omega_{i}}]+m_{i}\mathbb{E}[Y\mathds{1}_{\Omega_{i}}]$
$=\int_{B_{i}}Q^{-}_{X}(p)\mathrm{~{}d}\hat{h}(p)+\int_{B_{i}}Q^{-}_{Y}(p)\mathrm{~{}d}\hat{h}(p).$
For an interior point $p$ of $K$, Theorem 1 implies that $Q^{-}_{X+Y}(p)=Q^{-}_{X}(p)+Q^{-}_{Y}(P)$. Combining the above arguments, $I_{h_{c}}$ is $K$-additive.
Next suppose for some $p\in[0,1]$, $\mathrm{VaR}_{p}^{-}$ is not $K$-additive. By Theorem 1, $p\in(a_{i},b_{i}]$ for some $B_{i}=(a_{i},b_{i})$. Since $\hat{h}$ is continuous on $[a_{i},b_{i}]$, $\hat{h}_{l}$ is continuous at $p$ and $I_{h_{l}}$ does not have a mass on $\mathrm{VaR}_{p}^{-}$, this shows that $I_{h_{l}}$ is $K$-additive. A similar argument shows that $I_{h_{r}}$ is also $K$-additive.
For the only if direction, we construct counterexamples taking $\Omega=[0,1]$.
Case 1: If $\hat{h}$ is not linear on some $B_{i}=(a,b)$. Pick $X,Y$ such that $X=Y=0$ on $[0,a]$ and $X=Y=3$ on $[b,1]$. Since $I_{h}$ is not a multiple of the expectation on $(a,b)$, we can pick $X,Y$ to have range $[1,2]$ on $(a,b)$ so that $I_{h}$ restricted to $B_{i}$ is not additive. This can be done because $K$-concentration puts no restriction on the dependence structure of ($X|_{(a,b)},Y|_{(a,b)})$. Since $X=Y$ outside $(a,b)$, $I_{h}$ is additive when restricted to $(a,b)^{c}$. Overall, $I_{h}$ is not additive for $(X,Y)$.
Case 2: Suppoes for some $B_{i}=[a_{i},b_{i}]$, $\hat{h}$ is not right continuous at $a$. Then $I_{h}$ puts a mass on $Q^{-}_{a}$ and for some $\epsilon\in(0,a)$, $(a-\epsilon,a)\cap K=\emptyset$. A counterexample is illustrated by Figure 1. Let $U$ be a uniform $[0,1]$ random variable, $X=f(U)$ and $Y=g(U)$ with $f,g$ being the two functions in the illustration.
Case 3: A similar counterexample can be constructed where additivity fails due to a mass on some $Q^{+}_{b_{i}}$. ∎
[…]
Figure 1: An example where $Q_{a}^{-}$ and $\rho$ are not additive
Remark 2.
The following relationships are obvious from the definitions. Let $K_{1}\subseteq K_{2}$ be closed subsets of $[0,1]$.
  1. (a)
    For any random vector, $K_{2}$-concentration implies $K_{1}$-concentration.
  2. (b)
    For any distortion riskmetric, $K_{1}$-additivity implies $K_{2}$-additivity.
4 Function-indexed partial comonotonicity
For random variables $X,Y,Z$, if $(X,Z)$ and $(Y,Z)$ are comonotonic, it is in general not true that $(X,Y)$ is also comonotonic. That is, the property of having perfect comovements is not transitive. However, depending on $Z$, the property is partially transitive.
In this section, we define and investigate partial comonotonicity indexed by functions, and establish an equivalence to set-indexed comonotonicity. Let $\mathcal{G}$ be the collection of left-continuous increasing functions from $[0,1]$ to $\mathbb{R}$.
Definition 3.
Let $g\in\mathcal{G}$. A random vector $(X_{1},\dots,X_{d})$ is $g$-comonotonic if there exists a random variable $Z$ such that
  1. (a)
    the left quantile function $Q^{-}_{Z}$ is equal to $g$, and
  2. (b)
    $(X_{i},Z)$ is comonotonic for all $i=1,\dots,d$.
Definition 4.
Given $g\in\mathcal{G}$, a riskmetric $\rho$ is $g$-additive if for all random vector $(X_{1},\dots,X_{d})$ satisfying $g$-comonotonicity, we have $\rho(X_{1}+\cdots+X_{d})=\rho(X_{1})+\cdots+\rho(X_{d})$.
For $f,g\in\mathcal{G}$, we write $f\precsim g$ when there exists an increasing function $h$ such that $f=h\circ g$.
Proposition 4.
If $f\precsim g$, then $g$-comonotonicity implies $f$-comonotonicity, and in turn, $f$-additivity implies $g$-additivity.
Proof.
Let $h$ be such that $f=h\circ g$, and $Z$ be as in the definition of $g$-comonotonicity. Then $h(Z)$ satisfies the definition of $f$-comonotonicity. The other claim follows trivially. ∎
Remark 3.
[…]
  1. (a)
    $f$-comonotonicity is equivalent to $g$-comonotonicity;
  2. (b)
    $f\cong g$;
  3. (c)
    for all $0\leq a<b\leq 1$, $f(a)=f(b)$ if and only if $g(a)=g(b)$.
To establish a connection between the two notions of partial comonotonicity, we first define two functions.
Define the map $\mathcal{I}:\mathcal{G}\mapsto\mathcal{S}$ as follows. Given $g\in\mathcal{G}$, define an equivalence relation $\sim$ on $[0,1]$ by $a\sim b$ if and only if $g$ is constant on $[\min(a,b),\max(a,b)]$. Each equivalence class is an interval. Let $\{D_{\lambda}\}_{\lambda\in\Lambda}$ be the set of equivalence classes. Replace each $D_{\lambda}$ by its interior, and there are only countably many non-empty sets remaining, denoted by $(B_{i})_{i=1}^{\infty}$. Take $B=\bigcup_{i=1}^{\infty}B_{i}$ and $\mathcal{I}(g)$ to be $B^{c}$.
We next define the map $\mathcal{V}:\mathcal{S}\mapsto\mathcal{G}$. Given $K\in\mathcal{S}$, let $K_{i}=[c_{i},d_{i}]$ be the disjoint closed partition of $K$ and $B_{i}=(a_{i},b_{i})$ be the disjoint open partition of $K^{c}$. If $p\in(a_{i},b_{i})$ for some $i$, take $g(p)=a_{i}$. If $p\in[c_{i},d_{i}]$ for some $i$, take $g(p)=p$. Modify any discontinuities of $g$ such that $g$ is left-continuous. Define $\mathcal{V}(K)=g$. It is straightforward to check that $g$ is increasing.
Intuitively, $\mathcal{I}(g)$ is the set on which $g$ is strictly increasing, and $\mathcal{V}(K)$ is a function with slope $1$ on $K$, slope 00 on $K^{c}$.
Proposition 5.
The maps $\mathcal{V}$ and $\mathcal{I}$ have the following properties:
  1. (a)
    $\mathcal{V}(\mathcal{I}(g))\cong g$ for all $g\in\mathcal{G}$;
  2. (b)
    $\mathcal{I}(\mathcal{V}(K))=K$ for all $K\in\mathcal{S}$.
  3. (c)
    $g_{1}\precsim g_{2}\implies\mathcal{I}(g_{1})\subseteq\mathcal{I}(g_{2})$ for all $g_{1},g_{2}\in\mathcal{G}$;
  4. (d)
    $K_{1}\subseteq K_{2}\implies\mathcal{V}(K_{1})\precsim\mathcal{V}(K_{2})$ for all $K_{1},K_{2}\in\mathcal{S}$.
Proof.
The proofs are straightforward by the constructions of $\mathcal{I}$ and $\mathcal{V}$.
  1. (a)
    Let $f=\mathcal{V}(\mathcal{I}(g))$. Take $0\leq a<b\leq 1$, we show that $f(a)=f(b)$ if and only if $g(a)=g(b)$. If $g(a)=g(b)$, then $(a,b)\subseteq B_{i}=(a_{i},b_{i})$ for some $i$. By construction $f(a)=a_{i}=f(b)$. If $g(a)\neq g(b)$, let $B_{i}=(a_{i},b_{i})$ be the equivalence class such that $[a_{i},b_{i})$ contains $a$. Then $f(a)=a_{i}<b_{i}\leq f(b)$.
  2. (b)
    Let $L=\mathcal{I}(\mathcal{V}(K))$. Let $(a,b)$ be a maximal open interval contained in $K^{c}$ and $g=\mathcal{V}(K)$, we have $g\equiv a$ on $[a,b)$, and $g(x)\neq a$ for any $x\notin[a,b)$. Hence by construction, $(a,b)\cap L=\emptyset$ and $a,b\in L$, so that $(a,b)$ is also a maximal open interval contained in $L^{c}$.
  3. (c)
    When $g_{1}\precsim g_{2}$, for any $a,b\in[0,1]$ we have $g_{2}(a)=g_{2}(b)\implies g_{1}(a)=g_{1}(b)$. If $(a,b)$ is one of the $B_{i}$ in the construction of $\mathcal{I}(g_{2})$, then some $(a^{\prime},b^{\prime})$ with $a^{\prime}\leq a$ and $b\leq b^{\prime}$ is one of the $B_{i}$ in the construction of $\mathcal{I}(g_{1})$, so that $\mathcal{I}(g_{2})^{c}\subseteq\mathcal{I}(g_{1})^{c}$.
  4. (d)
    Let $g_{1}=\mathcal{V}(K_{1})$ and $g_{2}=\mathcal{V}(K_{2})$. Suppose for some $a,b\in[0,1]$ we have $g_{2}(a)=g_{2}(b)$, we show that $g_{1}(a)=g_{1}(b)$. By the definition of $\mathcal{V}$, $(a,b)\subseteq K_{2}^{c}$. Hence $(a,b)\subseteq K_{1}^{c}$ and $g_{1}(a)=g_{1}(b)$. ∎
The following theorem establishes the equivalence between the two notions of partial comonotonicity.
Theorem 3.
For $K\in\mathcal{S}$ and $g\in\mathcal{G}$, we have the following equivalences:
  1. (a)
    $K$-concentration is equivalent to $\mathcal{V}(K)$-comonotonicity;
  2. (b)
    $g$-comonotonicity is equivalent to $\mathcal{I}(g)$-concentration.
Proof.
(a) $\Rightarrow$: Let $g=\mathcal{V}(K)$ and suppose $\mathbf{X}$ is $K$-concentratated. Let $\{A_{p}\}_{p\in K}$ be a nested collection of common tail events, that is, $p>q\implies A_{p}\subseteq A_{q}$. Let $K_{i}=[c_{i},d_{i}]$ be the closed partition of $K$ and $B_{i}=(a_{i},b_{i})$ be the open partition of $K^{c}$. For $\omega\in\Omega$ let $p=\sup\{q:q\in K,\omega\in A_{q}\}$. Define $Z(\omega)$ according to the following rules.
  1. (i)
    If $\omega\notin A_{p}$, take $Z(\omega)=p$.
  2. (ii)
    If $\omega\in A_{p}$ and there exists a strictly decreasing sequence $p_{n}\downarrow p$ with $p_{n}\in K$ for each $n$, also take $Z(\omega)=p$.
  3. (iii)
    Otherwise, let $r=\inf\{q:q>p,q\in K\}$ and take $Z(\omega)=(r+p)/2$.
We first show that each $X_{i}$ is comonotonic with $Z$. Fix $i$, and take $\omega_{1},\omega_{2}\in\Omega$. Without loss of generality, assume $p_{1}=Z(\omega_{1})<Z(\omega_{2})=p_{2}$. If $p_{1}$ arises from case (i), pick $p=p_{1}$. If $p_{1}$ arises from case (ii), pick $p$ with $p_{1}<p<p_{2}$. If $p_{1}$ arises from case (iii), pick $p=r$. In all cases, $\omega_{2}\in A_{p}$ but $\omega_{1}\notin A_{p}$ and hence $X_{i}(\omega_{2})\geq X_{i}(\omega_{1})$.
(a)$\Leftarrow$: Let $\mathcal{V}(K)=g$. Suppose $\mathbf{X}$ is $g$-comonotonic. Let $Z$ be the random variable comonotonic with each $X_{i}$ with $Q_{Z}=g$. For each $p\in K$, define $A_{p}=\{\omega:Z(\omega)>p\}$. By the construction of $\mathcal{V}(K)$, $\mathbb{P}(A_{p})=1-p$ and for any $\omega\in A_{p},\omega^{\prime}\in A_{p}^{c}$, $Z(\omega)>Z(\omega^{\prime})$. Comonotonicity then requires $X_{i}(\omega)\geq X_{i}(\omega^{\prime})$ for each $i$, so that $\mathcal{X}$ is $p$-concentrated for each $p\in K$.
(b): Combining (a), Remark 3 and Proposition 5, we have the following equivalences:
$g$-comonotonicity $\Leftrightarrow\mathcal{V}(\mathcal{I}(g))$-comonotonicity $\Leftrightarrow\mathcal{I}(g)$-comonotonicity.∎
Remark 4.
Let $K\in\mathcal{S}$. Using the terminology of copulas, $K$-concentration represents a collection of copulas $\mathcal{C}_{K}$, in the sense that a random vector $\mathbf{X}$ is $K$-concentrated if and only if the copula of $\mathbf{X}$ is a member of $\mathcal{C}_{K}$. The best tool to describe $\mathcal{C}_{K}$ is ordinal sums from Nelson (2006, Section 3.2.2). Let the open intervals $\{B_{i}:=(a_{i},b_{i})\}$ be a disjoint partition of $K^{c}$ and closed intervals $\mathcal{K}=\{K_{i}:=(c_{i},d_{i})\}$ be a disjoint partition of $K$. Replace each $B_{i}$ by its closure $[a_{i},b_{i}]$. Let $\{J_{i}\}$ be the collection of non-degenerate closed intervals among all $\overline{B_{i}}$ and $K_{i}$. Any collection of copulas $\{C_{i}\}$ with the same index of $\{J_{i}\}$ defines a copula through the ordinal sum. Such an ordinal sum copula is a member of $\mathcal{C}_{K}$ if the following holds:
each $C_{i}$ is the comonotonic copula when $J_{i}\in\mathcal{K}$.
Conversely, all members of $\mathcal{C}_{K}$ is of this form. Figure 2 illustrates this idea. Take $K=[0,1/3]\cup[2/3,1]$. Put any copula inside the middle box, the resulting ordinal sum copula is in $\mathcal{C}_{K}$.
[…]
Figure 2: $K$-concentration as a collection of Ordinal sums
5 On spectral risk measures
So far, our study has not imposed monotonicity on the riskmetrics $I_{h}$ for $h\in\mathcal{H}^{\rm bv}$. Next, we consider a special class of $I_{h}$ that is monotone and convex,  [^1] We also assume that $h$ is continuous at 00. For a concave increasing function $h$ with $h(0)=0$ and $h(1)=1$, the only possible continuity is at 00. A jump corresponds to adding a multiple of the essential supremum. that is, the class of the spectral risk measures. We obtain an equivalent condition on the random vector for which a given spectral risk measure is additive.
Spectral risk measures, as introduced in Acerbi (2002), take the form
$\rho_{g}(X)=\int_{0}^{1}g(t)Q_{X}(t)\mathrm{~{}d}t$
for some increasing function $g:[0,1]\mapsto\mathbb{R}^{+}$ that integrates to $1$. This is a special case of a distortion riskmetric. Define a distortion riskmetric $I_{h}$ by letting the distortion function $h$ has its conjugate function $\hat{h}$ to be the integral of $g$, we have that $\rho_{g}=I_{h}$. In the special setting of spectral risk measures, we will be able to provide a result stronger than Theorem 2. We first rephrase Theorem 2 in this context.
Corollary 1.
Let $g_{1},g_{2}\in\mathcal{G}$. The spectral risk measure $\rho_{g_{1}}$ is $g_{2}$-additive if and only if $g_{1}\precsim g_{2}$.
Proof.
Let $I_{h}$ be the Choquet integral such that $I_{h}=\rho_{g}$. That is, $\hat{h}^{\prime}(t)=g(t)$. By Theorem 2, for $K\in\mathcal{S}$, $\rho_{g}$ is $K$-additive if and only if $\hat{h}$ is linear on each connected component of $K^{c}$, which is equivalent to $g$ being constant on each connected component of $K^{c}$. But $g$ is constant on each connected component of $\mathcal{I}(g)^{c}$, therefore $\rho_{g}$ is $K$-additive if and only if $K^{c}\subseteq\mathcal{I}(g)^{c}$. Hence $\rho_{g_{1}}$ is $g_{2}$-additive if and only if $\mathcal{I}(g_{1})\subseteq\mathcal{I}(g_{2})$, if and only if $g_{1}\precsim g_{2}$. ∎
Theorem 4.
Let $g\in\mathcal{G}$ and $\mathbf{X}$ be a random vector. The spectral risk measure $\rho_{g}$ is additive on $\mathbf{X}$ if and only if $\mathbf{X}$ is $g$-comonotonic.
The proof uses the following lemma.
Lemma 1.
(McNeil et al. (2015), Theorem 7.28) Let $X^{c},Y^{c}$ be comonotonic copies of $X,Y$. Then $\mathbb{E}(XY)\leq\mathbb{E}(X^{c}Y^{c})$ with equality holding if and only if $X,Y$ are comonotonic.
Proof of Theorem 4.
Let $Z$ be some random variable with $Q_{Z}(t)=g(t)$. Given a random vector $\mathbf{X}=(X,Y)$, consider the following expressions:
$\rho_{g}(X+Y)$ $=\int_{0}^{1}Q_{Z}(t)Q_{X+Y}(t)\mathrm{~{}d}t$
$\geq\mathbb{E}\big{(}Z(X+Y)\big{)}=\mathbb{E}(ZX)+\mathbb{E}(ZY),$
and
$\mathbb{E}(ZX)+\mathbb{E}(ZY)$ $\leq\int_{0}^{1}Q_{Z}(t)Q_{X}(t)\mathrm{~{}d}t+\int_{0}^{1}Q_{Z}(t)Q_{Y}(t)\mathrm{~{}d}t$
$=\rho(X)+\rho(Y).$
()::absent($\Leftarrow$):( ⇐ ) : If $(X,Y)$ is $g$-comonotonic, pick $Z$ as in the definition so that $(X,Z)$ and $(Y,Z)$ are comonotonic. Hence, $Z$ and $X+Y$ are comonotonic, so all inequalities become equalities.
()::absent($\Rightarrow$):( ⇒ ) : Conversely, let $U$ be the distributional transform (Föllmer and Schied (2016, Lemma A.32)) of $X+Y$, i.e., $X+Y=Q_{X+Y}(U)$. Let $Z=g(U)$, $Z$ and $X+Y$ are comonotonic because they are both increasing functions of $U$. Then the first inequality is an equality, but $Z$ cannot be comonotonic with both $X$ and $Y$, so the second inequality is strict. ∎
Remark 5.
Theorem 4 is stronger in the following sense. In Theorem 2, an $\mathbf{X}$ not satisfying the corresponding $K$-concentration may still be additive under $I_{h}$: it only requires that any $\mathbf{X}$ satisfying $K$-concentration be additive, and that some $\mathbf{X}$ not satisfying $K$-concentration be not additive. However, in the context of spectral risk measures, any $\mathbf{X}$ not satisfying $K$-concentration cannot be additive.
Example 2.
For $p\in(0,1)$, let $g_{p}(x)=\mathds{1}_{[p,1]}/(1-p)$ on $x\in[0,1]$. The spectral risk measure induced by $g_{p}$ is $\mathrm{ES}_{p}$. The corresponding closed set $K$ is $\mathcal{I}(g_{p})=\{p\}$. Theorem 2 would imply that $\mathrm{ES}_{p}$ is $p$-additive, and that any $p$-additive distortion riskmetric $\rho$ is of the form $\rho=\alpha\mathrm{ES}_{p}+\beta\mathbb{E}$ for some $\alpha,\beta\in\mathbb{R}$. Theorem 4 would further imply that, if $\mathbf{X}$ is not $p$-concentrated, then $\mathrm{ES}_{p}$ is not additive on $\mathbf{X}$.
6 Conclusion
We provided two equivalent formulation of partial comonotonicity: $K$-concentration, indexed by some closed set $K\subseteq[0,1]$ and $g$-comonotonicity, with $g:[0,1]\mapsto\mathbb{R}$ an increasing left-continuous function. For a dependence structure $D$ from the above class, we characterized the class of distortion riskmetrics that are additive for random vectors satisfying $D$.
In place of additivity, one may consider another interesting property: aversion, where risks are highest under a certain dependence structure $D$. Formally, a riskmetric $\rho$ is $D$-averse if for any $(X,Y)$ satisfying $D$, any $X^{\prime}\buildrel\mathrm{d}\over{=}X$ and $Y^{\prime}\buildrel\mathrm{d}\over{=}Y$ satisfies
$\rho(X+Y)\geq\rho(X^{\prime}+Y^{\prime}).$
Under the assumption of subadditivity for the riskmetric, $D$-additivity implies $D$-aversion. The validity of the converse condition reduces to whether additivity is always achievable, which is an interesting problem in itself.
When $D$ is comonotonicity, $D$-aversion was studied in Mao and Wang (2020) and termed diversification consistency. When $D$ is $p$-concentration, $D$-aversion was studied in Han et al. (2024), and characterization results were established. For a generic instance $D$ of partial comonotonicity, the analysis seems more challenging, and a theory remains to be established.